- Emergency Help
- Computer Training for All Staff
- Blog Policy and Procedure
- Middle Managers Meeting Minutes
George A. Smathers Libraries
University of Florida
2124 NE Waldo Rd
Gainesville, FL 32609
Telephone: (352) 273-2800
Fax: (352) 392-2354
Information Technology Staff Meeting Notes
April 19, 2011
April 19, 2011
Present: Shruti Bagri, Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Michael Jay, Tom Minton, Chris Nicolich, Amy Polk, Patricia Ruwell, Josh Spurgin, Mike Tyler, Evan WackAnnouncements
Ongoing IT Updates
- Spring Picnic April 29.
- TSS - April 13 highlights:
Times New Roman 12 is the official Library font. Time to start gathering milestones accomplished in the past year.
Interim Storage Plan - new High Density storage pushed back to 2015/16. The Library is currently looking for an interim storage location to house statewide collection. DLC/Preservation will be relocated to the interim location.
- Annual TSS meeting scheduled for July 19. Will include a Project Management exercise in which the winning project will be funded. Will require loaner laptops to be reserved and configured with Project Management templates.
- Clean Slate deadline - no word. Last build had only 2 problems.
- Altiris - Asset Management - purging database to let Altiris rebuild it. New SQL instance is complete. Trish and Evan rebuilt image deployment jobs and are creating final images. Staff image is done. Will set up machines in West next week for staff to come and test their departmental software.
- Technology room guides - East conference and 1A still in progress. 1A is 75% complete.
- Public computer cable cleanup in progress. Need to purchase miscellaneous cables to complete the project.
- Staff computers - list verified, sending out departmental emails to advise users.
- Archivist Toolkit - Logan to do this week.
- East conference room - we'd like to reconfigure it, mounting the computer under the table and getting rid of the podium. Will to bring up the topic at TSS.
- IT SharePoint deployment complete. 10 machines did not get it - Ref Desk computers - had to install remotely.
- Departmental printing mapping - Josh is working on it.
- Web unit head search - 4 candidates. Phone interviews to be completed by Friday.
- Testing new version of Privilege Authority.
- Testing new version of EPO.
- Checkout laptop changes - auto VPN connection, managed through Group Policy, using Bit Locker. Will use restore points to reset the machines once they are returned. Sounds good, must thoroughly test and demo before implementing.
- West Group Study rooms - awaiting installation date.
- Education mobile video cart - PO received, awaiting installation date.
- MSL107 GIS/3D modeling - should get word next week.
- Help Desk - finals are next week, and then break week and then Summer A starts - May 9. Kelly is leaving for the summer, but Diana thinks she is getting Work Study and can continue working. Debra submitted a requisition to post an ad for new students for summer semester to HR today.
DLC General Issues
- Recruitment Study - Winston has a phone call with Brian at 3:30PM today. Michael Jay is getting set up with Chris ready to assist.
- Document Disposition database - several weeks left. Phase II demoed. Over 4,000 entries have been submitted from around the state.
- Kuale OLE - phone interviews this week - 9 applicants.
- ITS - changes to be demoed this afternoon to Missy and Isabel.
- Class Registration System from HSCL - Chris is reviewing.
- Library Hours - one week of work left.
- Civil War microfilm database - it is being donated to the Library (DLC) need to figure out how to tag the data structure.
- Mark is working on the migration to Solr/Lucene.
- SharePoint training class yesterday - went fairly well. Amy will be teaching at MSL later this week.
- VIVO - no word. Both old and new servers are online. Meeting with VIVO and Admin on April 21.
- MS Forefront - 90% done.
- Server Room - all surveyed stuff accounted for, will send list to Bill Hanssen. Tentatively scheduled to move to colocation spot in two weeks. Need to figure out what to do with EPO, app development, etc. Will need to move.
- DHCP - ready to go. Will start on small branches tomorrow.
- CNS infrastructure - ballpark price for interim storage locations.
- Ticket logged with CNS for wireless repeaters in Room 1A.
- People Counter - IP assignment in process. Vendor to install.
- New PHP/MySQL/Tomcat server - WIN2K8 R2 based. Ease management for existing staff. Will replace the existing Linux server.
- Cynthia is confused on where ALEPH draws its Cataloger info.
- Study Room Reservations - no one took into account the effect of extended hours on study room scheduling. We are working on the issue, unsure if it will be resolved before finals. The upgrade to IE9 on the computer has helped with overall performance.
- Catalog log-in shift to Gatorlink - added an extra button for Gatorlink log-on. Issue still exists with Emeritus group.
- Meeting for app issues rescheduled for June.
- FEFDL - ongoing.